OpenStudy (anonymous):

I think I can solve this problem, but not entirely sure. Help? It's absolute value.

OpenStudy (anonymous):

Solve for x: −3|2x + 6| = −12 x = 1 and x = 5 x = −1 and x = −5 x = −9 and x = 3 No solution

OpenStudy (anonymous):

@amistre64 Could you show me?

OpenStudy (anonymous):

@Nnesha ?

Nnesha (nnesha):

yes first we need to get rid of everything at left side except the absolute bar so move -3 to the right side

OpenStudy (anonymous):

Divide?

Nnesha (nnesha):

right divide both sides by -3

Nnesha (nnesha):

let me know what you get after that so we can continue :=))

OpenStudy (anonymous):

um. -4

Nnesha (nnesha):

no negative divide by negative = positive

Nnesha (nnesha):

-12/-3= ?

OpenStudy (anonymous):

Oh, okay. So +4

OpenStudy (anonymous):

Positive 4.

Nnesha (nnesha):

right \[\left| 2x+ 6\right|=4\] absolute contains two solution \[\left| 2x+6 \right| \rightarrow (2x+6)^2\] absolute bar of 2x+6 is same as square of (2x+6) and when take square root we should get plus/minus answer \[\sqrt{x^2}= \pm x\]

OpenStudy (anonymous):

What?

Nnesha (nnesha):

\[\left| 2x+6 \right|=4 \rightarrow 2x+6=4 ~\rm and~ 2x+6= -4\] now solve both equation for x

OpenStudy (anonymous):

Subtract 6 from both sides?

Nnesha (nnesha):

if the absolute bar equal to positive number then you will get 2 solution here is an example \[\left| a+b \right|=c\]\[a+b = c ~~\rm and ~~ a+b=-c\]

Nnesha (nnesha):

yes right

OpenStudy (anonymous):

Okay so 2x=-2?

Nnesha (nnesha):

for first equation yes 2x = -2

Nnesha (nnesha):

solve for x

OpenStudy (anonymous):

Divide. x= -1?

Nnesha (nnesha):

right

Nnesha (nnesha):

2nd equation

OpenStudy (anonymous):

-6 both sides?

Nnesha (nnesha):

right

OpenStudy (anonymous):

So, -10.

OpenStudy (anonymous):

2x=-10

OpenStudy (anonymous):

-5?

Nnesha (nnesha):

looks good

OpenStudy (anonymous):

So, it's b?

Nnesha (nnesha):

i think so.

OpenStudy (anonymous):

Okay. Thank you! :)

Nnesha (nnesha):

np :=))